# Classifying Infinity Topoi via weighted limits

### v1

@@ 3. Internal Higher Category Theory in an $\infty$-topos
$E_{1}=\Delta^{0}\coprod_{\Delta^{\{0,2\}}}\Delta^{3}\coprod_{\Delta^{\{1,3\}}}\Delta^{0}$
. This is known as the walking equivalence.
@@
induced by the spine inclusion $I_{n}$ and the map $E^{1}\rightarrow\Delta^{0}$ are equivalences of $\infty$-categories. We will denote by $\mathrm{Cat}(\mathcal{B})$ the full subcategory of $\mathcal{B}_{\Delta}$, whose objects are the $\mathcal{B}$-categories.
@@
In [Mar22, Definition 3.1.5 and Definition 3.2.1], a $\mathcal{B}$-category is defined as an object of $\mathcal{B}_{\Delta}$ which is internally local to the maps $I_{n}:$ and $E^{1}\rightarrow\Delta^{0}$, and [Mar22, Propositoin 3.2.7] shows that this definition is equivalent to the one given above.
@@
$Cat(\mathcal{A})\leftrightarrows Cat(\mathcal{B})$
@@
$const_{\mathcal{B}}:Cat_{\infty}\leftrightarrows Cat(\mathcal{B}):\Gamma_{\mathcal{B}}$
@@
    (Cotensoring) $(-)^{(-)}:=(const_{\mathcal{B}}(-),(-))$
@@
We will write $CAT_{\mathcal{B}}$ for the $(\infty,2)$-category from [Mar22, Remark 3.4.3], whose objects are $\mathcal{B}$-categories and such that $Hom_{CAT_{\mathcal{B}}}(X,Y)=Fun_{\mathcal{B}}(X,Y)$.
@@
${Hom_{C}(c,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d_{0},d_{1}}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(c,d)}$${C_{0}\times C_{0}}$
@@
$\Omega(A)\simeq\mathcal{B}_{/A}$
Given two objects of $\mathcal{B}_{/A}$, which can be viewed as objects of $\Omega$ in context $A$, we have an equality
$map_{\Omega}(G,H)\simeq Fun_{\mathcal{B}_{/A}}(G,H)$
For a $\mathcal{B}$-category $C$, we write $\underline{Psh}_{\mathcal{B}}(C):=\underline{Fun}_{\mathcal{B}}(C,\Omega_{\mathcal{B}})$
@@
We say that a large $\mathcal{B}$-category is presentable if it is equivalent to a $\mathcal{B}$-categorical localization of $\underline{Psh}_{\mathcal{B}}(C)$ for some small $\mathcal{B}$-category $C$ (see [MW25, Section 2.4] for more details) We say that a $\mathcal{B}$-category $C$ is a $\mathcal{B}$-topos if it satisfies the following conditions:
@@
    It satisfies descent. That is, the slice functor $C_{/-}:C\rightarrow Cat_{\mathcal{B}}$ is cocontinuous (see [MW25, Section 3.1] for more details)
@@
In [Joh02], the morphisms in the 2-category of $\mathcal{B}$-topoi are defined to be geometric morphisms, as opposed to algebraic morphisms. Thus, $\mathcal{TOP}_{\mathcal{B}}$ is an $\infty$-categorical analogue of the opposite of the 2-category of $\mathcal{B}$-topoi defined in [Joh02]
@@ 4. A Lemma on Flat Functors
$Filt((C)^{\Delta^{1}},\mathcal{S})\simeq Filt((C),\mathcal{S})^{\Delta^{1}}$
@@
Suppose that $S:C^{op}\rightarrow\mathcal{S}$ is a filtered functor of $\infty$-groupoids. Then the following are true for $R=\pi_{0}S^{K}$.
@@
    There exists some $c\in C$ such that $S(c)$ is nonempty
@@
    For any $x_{i}\in R(c_{i}),i=1,2$, there exists morphisms $f_{i}:e\rightarrow c_{i},i=1,2$ and $x_{3}\in R(e)$ such that $R(f_{i})(x_{3})=x_{i}i=1,2$
@@
    Suppose that we have morphisms $f_{i}:d\rightarrow c$ in $\mathcal{X}$ and $x\in X(d)$. Then there exists a $g:e\rightarrow d$ and $y\in X(e)$ such that $g\circ f_{1}=g\circ f_{2}$ and $g(y)=x$.
In particular $\pi_{0}(S)$ is a filtered functor of sets in the sense of [Joh02]
@@
$(K\coprod K)^{\triangleleft}\rightarrow D$
@@
Now, choose equivalences $\phi_{i}:S(c_{i})\simeq\mathbf{Un}(S)^{-1}(c_{i})$, $(y_{1},y_{2})=\phi((K\coprod K)\times\{0\})$. Let $f_{i}=Un(S)\circ h_{i}(\{n\}\times\Delta^{1})$ . By the definition of unstraightening, we have $S(f_{i})(e)$ is equivalent in $S(c_{i})$ to $y_{i}$. In particular, $\pi_{0}S^{K}(f_{i})(e_{i})=y_{i}$.
@@
where the natural transformation is induced by the evaluation map $ev:\Delta^{1}\times X^{\Delta^{1}}\rightarrow X$.
@@
suppose that $F:I\rightarrow Cat_{\infty}$ is a diagram. Let $I_{\neq\emptyset}$ be the full subcategory of $I$ consisting of objects $i$ such that $F(i)\neq\emptyset$. The $colimF=colimiF|_{I_{\neq\emptyset}}$
@@
$T_{i}(Hom_{C^{\Delta^{1}}}(-,y))(x)=colim_{C^{d^{i}}(z)=y}Hom_{C^{\Delta^{1}}}(-,z))=Hom_{C^{\Delta^{1}}}(x\circ C^{s^{i}},y)$
@@
By the characterization of filtered functors from [Lur09a, Section 5.3.5], it suffices to show that $\mathfrak{S}$ takes transformations of filtered colimits of representables to filtered colimits of representables. Since every map of ind objects has a level representation (reference?? this is true classically), we reduce to showing that $\mathfrak{S}$ sends diagrams of representables to representables.
Consider a natural transformation $\alpha:Hom_{C}(-,x)\xrightarrow{Hom_{C}(-,q)}Hom_{C}(-,y)$ For a morphism $f:z\rightarrow w$ in $C^{\Delta^{1}}$, the description of mapping spaces in functor categories as ends from [Hau22] implies that we can describe $Map_{C^{\Delta^{1}}}(f,q)$ as the limit of
@@
By the proof of Lemma 4.4, this can be identified with a map $\Phi$ which takes $z_{1}\in U(f)$ to the element of the coend represented by $(id_{dom(f)},z_{1})$. Similarly, we can construct a map $\Psi:U\rightarrow T_{1}(U)$ which takes $z_{1}\in U(f)$ to the element of the coend represented by $(id_{dom(f)},z_{1})$ We let the $\epsilon:U\rightarrow\mathfrak{S}\mathbf{T}U$ be the natural transformation induced by the universal property of pullback in the commutative diagram:
@@
We want to show that $\epsilon_{X}$ is an equivalence for $X\in Filt((C^{op})^{\Delta^{1}},\mathcal{S})$. By the $\infty$-categorical Yoneda lemma, it suffices to show that
$\pi_{0}\epsilon_{X}^{K}$
@@
$\epsilon_{X}^{K}=\epsilon_{X^{K}}$
and $\pi_{0}X^{K}$ is filtered by Lemma 4.3, it suffices to prove that $\pi_{0}\epsilon_{X}$ is an equivalence.
By applying $\pi_{0}$ to $\epsilon$ and corestricting to filtered functors, we obtain a natural transformation $\epsilon^{\prime}:\pi_{0}\rightarrow\pi_{0}\mathbb{S}\mathbb{T}$. Since $\epsilon^{\prime}$ is constructed identically to the counit of the equivalence from [Joh02, Proposition B.4.1.2], we can repeat the argument there and use Lemma 4.3. to show that $\mathbf{\epsilon}^{\prime}$ is a bijection.
@@
Let $p:Fun(\Delta^{1},C^{op})_{/id_{C^{op}}}\rightarrow Fun(\Delta^{1},C)$ be the projection map Then by Lemma 4.6 we can identify $\mathfrak{S}\mathbf{T}(U)_{0}(c)$ with the coend
$\int^{f}Hom_{C}(c,dom(p(f)))\times\mathfrak{S}\mathfrak{T}(U)(p(f))\big)\simeq\mathfrak{S}\mathbf{T}(U)(id_{c})\times hom(c,c)\simeq\mathfrak{S}(U)(id_{c})=\mathbf{T}_{0}(U)(c)$
where the two isomorphisms follow from the fact that $Fun_{\mathcal{B}}(\Delta^{1},\mathcal{X})_{/id_{c}},hom(c,c)$ each have a terminal object ($id_{c}$), and the final isomorphism is immediate from the definition of $\mathfrak{S}$.
We will define $\eta_{1}:\mathbb{T}\mathbb{S}(-)_{1}\rightarrow(-)_{1}$ be the formula:
$\int^{f}(id,\alpha_{dom(f)}p^{1}_{\alpha}):\mathbb{T}\mathbb{S}(U)_{1}(c)=\int^{f\in C^{\Delta^{1}}}Hom_{C}(c,codom(f))\times S(\alpha)(f)\\ \longrightarrow\int^{f\in C^{\Delta^{1}}}Hom_{C}(c,codom(f))\times U_{1}(f)\rightarrow U_{1}(c)$
where two right horizontal arrows are the equivalence of Lemma 4.3. Using Lemma 4.3, the map can be identified with the map given by
@@
we can reduce as before to showing that $\pi_{0}\eta_{1}$ is an equivalence. But since this unit component is constructed almost identically to the component of the unit of the equivalence from [Joh02, Proposition B.4.1.2], we can show $\pi_{0}\eta_{1}$ is an equivlaence usng Lemma 4.4 , Eq. 3, as well as the argument from [Joh02, Proposition B.4.1.2].
@@ 5. The $\infty$-Bicategory of Bounded Topoi
Suppose that $f^{*}:\mathcal{B}\rightarrow\mathcal{C}$ is geometric morphism, which induces a map. Then we have an equivalence $f^{*}(X^{K})\simeq f^{*}(X)^{K}$, natural in $\mathcal{B}$-categories $X$ and finite $\infty$-categories $K$
@@
whee the second equivalence comes from the universal property of $Psh(C)$ as the universal cocompletion. Under this equivalence, we see that $\Omega_{\mathcal{B}}$ can be identified with the presheaf $c\mapsto Psh(C)_{/y_{c}}\cong\mathcal{S}$. That is, it is the constant presheaf with value $\mathcal{S}$.
@@
$Fun_{\mathcal{B}}(X^{\Delta^{1}},Y)\simeq\int_{c}Fun(X(L(c))^{\Delta^{1}},Y(L(c)))\rightarrow\int_{c}Fun(X(L(c)),Y(L(d)))^{\Delta^{1}}\simeq\\ \int_{c}Fun(X(Lc),Y(Ld)^{\Delta^{1}})\simeq Fun_{\mathcal{B}}(X,Y^{\Delta^{1}})\simeq Fun_{\mathcal{B}}(X,Y)^{\Delta^{1}}$
@@
Note that this map is natural with respect to maps $Cat(\mathcal{B})\leftrightarrows Cat(\mathcal{A})$ induced by geometric morphisms of $\infty$-topoi
@@
Let $Fin_{\mathcal{B}}$ be the collection of $\mathcal{B}$-categories which can be identitied with locally constant functors of finite $\infty$-categories on $\mathcal{B}$. A $B$-category $J$ is said to be filtered if the colimit functor $colim:\underline{Fun}_{\mathcal{B}}(I,\Omega)\rightarrow\Omega$ is continuous for each $I\in Fin_{\mathcal{B}}$.
We say that a functor $f:C\rightarrow\Omega_{\mathcal{B}}$ is filtered if it is contained in the free $Fin_{\mathcal{B}}$ cocompletion of $C$
@@
$Filt_{\mathcal{B}}(X^{\Delta^{1}})\rightarrow Filt_{\mathcal{B}}(X)^{\Delta^{1}}$
@@
(6) ${Fun_{\mathcal{B}}^{cc}(\underline{Psh}_{\mathcal{B}}(C),\mathcal{X})^{\Delta^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{((h_{C})^{*})^{\Delta^{1}}}$${Fun_{\mathcal{B}}(C,\mathcal{X}))^{\Delta^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}C,\Omega_{\mathcal{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{B}}(f^{*}(C),\Omega_{\mathcal{X}})^{\Delta^{1}}}$${Fun_{\mathcal{B}}^{cc}(\underline{Psh}_{\mathcal{B}}(C^{\Delta^{1}}),\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{((h_{C})^{*})}$${Fun_{\mathcal{B}}(C^{\Delta^{1}},\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}(C^{\Delta^{1}}),\Omega_{\mathcal{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}(C)^{\Delta^{1}},\Omega_{\mathcal{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where $h_{C}$ is the internal version of the Yoneda imbedding from [MW25], and the vertical maps come from the natural map of 5.4. The diagram commutes by the naturality of the construction of 5.4, Lemma 5.2 as well as the identification $f_{*}\Omega_{\mathcal{X}}=\mathcal{X}$.
@@
In this case 5.4 reduces to proving that the natural map $Filt(C(c)^{\Delta^{1}},\mathcal{X}(c))\simeq Filt(C(c),\mathcal{X}(c))^{\Delta^{1}}$ from 4.5 is an equivalence. Since $X(c)$ is a full subcategory of $\mathcal{S}$ by Example 5.3, the result follows from Theorem 4.2
@@
By the dual of Corollary 2.7, it suffices to show that $TOP_{\mathcal{B}}$ has a tensoring with $\Delta^{1}$ and its underlying $\infty$-category has finite colimits. The second point, is [MW25, Propositoin 3.2.6.6]. For the first point, note that if $\mathcal{X}=Psh_{\mathcal{B}}(C)$, then $\mathcal{X}\otimes\Delta^{1}$ exists by Lemma 5.7. In general, choose an accessible localization $L:sPsh_{\mathcal{B}}(C)\leftrightarrows\mathcal{X}:i$. In this case, it is easy to see that if we form the pushout
@@
then $P$ satisfies the universal property of $\mathcal{X}\otimes\Delta^{1}$

### v2

@@ 3. Internal Higher Category Theory in an $\infty$-topos
$E_{1}=\Delta^{0}\coprod_{\Delta^{\{0,2\}}}\Delta^{3}\coprod_{\Delta^{\{1,3\}}}\Delta^{0}.$
This is known as the walking equivalence.
@@
induced by the spine inclusion and the map $E^{1}\rightarrow\Delta^{0}$ are equivalences of $\infty$-categories. We will denote by $\mathrm{Cat}(\mathcal{B})$ the full subcategory of $\mathcal{B}_{\Delta}$, whose objects are the $\mathcal{B}$-categories.
@@
In [Mar22, Definition 3.1.5 and Definition 3.2.1], a $\mathcal{B}$-category is defined as an object of $\mathcal{B}_{\Delta}$ which is internally local to the maps $I_{n}$ and $E^{1}\rightarrow\Delta^{0}$, and [Mar22, Proposition 3.2.7] shows that this definition is equivalent to the one given above.
@@
$Cat(\mathcal{A})\leftrightarrows Cat(\mathcal{B}).$
@@
$const_{\mathcal{B}}:Cat_{\infty}\leftrightarrows Cat(\mathcal{B}):\Gamma_{\mathcal{B}}.$
@@
    (Cotensoring) $(-)^{(-)}:=Fun_{\mathcal{B}}(const_{\mathcal{B}}(-),(-))$
@@
We will write $CAT_{\mathcal{B}}$ for the $\infty$-bicategory from [Mar22, Remark 3.4.3], whose objects are $\mathcal{B}$-categories and such that $Hom_{CAT_{\mathcal{B}}}(X,Y)=Fun_{\mathcal{B}}(X,Y)$.
@@
${Hom_{C}(c,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(d_{0},d_{1})}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(c,d)}$${C_{0}\times C_{0}}$
@@
$\Omega(A)\simeq\mathcal{B}_{/A}.$
Given two objects of $\mathcal{B}_{/A}$, which can be viewed as objects of $\Omega$ in context $A$, we have an equivalence
$map_{\Omega}(G,H)\simeq Fun_{\mathcal{B}_{/A}}(G,H).$
For a $\mathcal{B}$-category $C$, we write $\underline{Psh}_{\mathcal{B}}(C):=\underline{Fun}_{\mathcal{B}}(C,\Omega_{\mathcal{B}})$.
@@
We say that a large $\mathcal{B}$-category is presentable if it is equivalent to a $\mathcal{B}$-categorical localization of $\underline{Psh}_{\mathcal{B}}(C)$ for some small $\mathcal{B}$-category $C$ (see [MW25, Section 2.4] for more details). We say that a $\mathcal{B}$-category $C$ is a $\mathcal{B}$-topos if it satisfies the following conditions:
@@
    It satisfies descent. That is, the slice functor $C_{/-}:C\rightarrow Cat_{\mathcal{B}}$ is cocontinuous (see [MW25, Section 3.1] for more details).
@@
In [Joh02], the morphisms in the 2-category of $\mathcal{B}$-topoi are defined to be geometric morphisms, as opposed to algebraic morphisms. Thus, $TOP_{\mathcal{B}}$ is an $\infty$-categorical analogue of the opposite of the 2-category of $\mathcal{B}$-topoi defined in [Joh02].
@@ 4. A Lemma on Flat Functors
$Filt((C)^{\Delta^{1}},\mathcal{S})\simeq Filt((C),\mathcal{S})^{\Delta^{1}}.$
@@
Suppose that $S:C^{op}\rightarrow\mathcal{S}$ is a filtered functor of $\infty$-categories. Then the following are true for $R=\pi_{0}S^{K}$:
@@
    There exists some $c\in C$ such that $S(c)$ is nonempty.
@@
    For any $x_{i}\in R(c_{i}),i=1,2$, there exists morphisms $f_{i}:e\rightarrow c_{i},i=1,2$ and $x_{3}\in R(e)$ such that $R(f_{i})(x_{3})=x_{i}\,\,i=1,2$.
@@
    Suppose that we have morphisms $f_{i}:d\rightarrow c$ in $C$ and $x\in X(d)$. Then there exists a $g:e\rightarrow d$ and $y\in X(e)$ such that $g\circ f_{1}=g\circ f_{2}$ and $g(y)=x$.
In particular, $\pi_{0}(S)$ is a filtered functor of sets in the sense of [Joh02].
@@
$(K\coprod K)^{\triangleright}\rightarrow D$
@@
Now, choose equivalences $\phi_{i}:S(c_{i})\simeq\mathbf{Un}(S)^{-1}(c_{i})$, $(y_{1},y_{2})=\phi|_{(K\coprod K)\times\{0\}}$. Let $f_{i}=Un(S)\circ H|_{(K\coprod K)\times\{i\}}$. By the definition of unstraightening, we have $S(f_{i})(e)$ is pointwise equivalent in $S(c_{i})^{K}$ to $y_{i}$. In particular, [Lur09a, Corollary 5.1.2.3] implies that $\pi_{0}S^{K}(f_{i})(e_{i})=y_{i}$.
@@
where the natural transformation is induced by the evaluation map $ev:\Delta^{1}\times C^{\Delta^{1}}\rightarrow C$.
@@
suppose that $F:I\rightarrow Cat_{\infty}$ is a diagram. Let $I_{\neq\emptyset}$ be the full subcategory of $I$ consisting of objects $i$ such that $F(i)\neq\emptyset$. The $colimF=colimF|_{I_{\neq\emptyset}}$
@@
$T_{i}(Hom_{C^{\Delta^{1}}}(-,y))(x)=colim_{C^{d^{i}}(z)=x}Hom_{C^{\Delta^{1}}}(x,z))=Hom_{C^{\Delta^{1}}}(x\circ C^{s^{i}},y)$
@@
By the characterization of filtered functors from [Lur09a, Section 5.3.5], it suffices to show that $\mathfrak{S}$ takes transformations of filtered colimits of representables to filtered colimits of representables. Since every map of ind objects has a level representation, we reduce to showing that $\mathfrak{S}$ sends diagrams of representables to representables.
Consider a natural transformation $\alpha:Hom_{C}(-,x)\xrightarrow{Hom_{C}(-,q)}Hom_{C}(-,y)$. For a morphism $f:z\rightarrow w$ in $C^{\Delta^{1}}$, the description of mapping spaces in functor categories as ends from [Hau22] implies that we can describe $Map_{C^{\Delta^{1}}}(f,q)$ as the limit of
@@
By the proof of Lemma 4.4, this can be identified with a map $\Phi$ which takes $z_{1}\in U(f)$ to the element of the coend represented by $(id_{dom(f)},z_{1})$. Similarly, we can construct a map $\Psi:U\rightarrow T_{1}(U)$ which takes $z_{1}\in U(f)$ to the element of the coend represented by $(id_{dom(f)},z_{1})$. We let the $\epsilon_{U}:U\rightarrow\mathfrak{S}\mathbf{T}U$ be the natural transformation given pointwise by the universal property of pullback in the commutative diagram:
@@
We want to show that $\epsilon_{U}$ is an equivalence for $U\in Filt((C^{op})^{\Delta^{1}},\mathcal{S})$. By the $\infty$-categorical Yoneda lemma, it suffices to show that
$\pi_{0}\epsilon_{U}^{K}$
@@
$\epsilon_{U}^{K}=\epsilon_{U^{K}}$
and $\pi_{0}U^{K}$ is filtered by Lemma 4.3, it suffices to prove that $\pi_{0}\epsilon_{U}$ is an equivalence. Since $\epsilon^{\prime}$ is constructed identically to the counit of the equivalence from [Joh02, Proposition B.4.1.2], we can repeat the argument there and use Lemma 4.3. to show that $\mathbf{\epsilon}^{\prime}$ is a bijection.
@@
Let $p:Fun(\Delta^{1},C^{op})_{/id_{C^{op}}}\rightarrow Fun(\Delta^{1},C)$ be the projection map. Then we can identify $\mathfrak{S}\mathbf{T}(U)_{0}(c)$ with the coend
$\int^{f}Hom_{C}(c,dom(p(f)))\times\mathfrak{S}\mathbf{T}(U)(p(f))\simeq\mathfrak{S}\mathbf{T}(U)(id_{c})\times hom(c,c)\simeq\mathfrak{S}(U)(id_{c})=\mathbf{T}_{0}(U)(c)$
where the first two isomorphisms follow from the fact that $Fun(\Delta^{1},C)_{/id_{c}},hom(c,c)$ each have a terminal object ($id_{c}$) as well as Lemma 4.6, and the final isomorphism is immediate from the definition of $\mathfrak{S}$.
Let $\alpha:U_{0}\rightarrow U_{1}$ be a natural transformation of functors $C\rightarrow D$. We will define $\eta_{1}:\mathbb{T}\mathbb{S}(\alpha)_{1}\rightarrow(\alpha)_{1}$ be the formula:
$\int^{f}(id,\alpha_{dom(f)}p^{1}_{\alpha}):\mathbb{T}\mathbb{S}(U)_{1}(c)=\int^{f\in C^{\Delta^{1}}}Hom_{C}(c,codom(f))\times\mathbb{S}(\alpha)(f)\\ \longrightarrow\int^{f\in C^{\Delta^{1}}}Hom_{C}(c,dom(f))\times U_{1}(f)\rightarrow U_{1}(c)$
where last horizontal arrow is the equivalence of Lemma 4.4. Using Lemma 4.4, the map can be identified with the map given by
@@
we can reduce as before to showing that $\pi_{0}\eta_{1}$ is an pointwise bijection. But since this unit component is constructed almost identically to the component of the unit of the equivalence from [Joh02, Proposition B.4.1.2], we can show $\pi_{0}\eta_{1}$ is an equivalence usng Lemma 4.3 , Eq. 3, as well as the argument from [Joh02, Proposition B.4.1.2].
@@ 5. The $\infty$-Bicategory of Bounded Topoi
Suppose that $f^{*}:\mathcal{B}\rightarrow\mathcal{C}$ is geometric morphism. Then we have an equivalence $f^{*}(X^{K})\simeq f^{*}(X)^{K}$, natural in $\mathcal{B}$-categories $X$ and finite $\infty$-categories $K$.
@@
whee the second equivalence comes from the universal property of $Psh(C)$ as the free cocompletion. Under this equivalence, we see that $\Omega_{\mathcal{B}}$ can be identified with the presheaf $c\mapsto Psh(C)_{/y_{c}}\cong\mathcal{S}$. That is, it is the constant presheaf with value $\mathcal{S}$.
@@
$Fun_{\mathcal{B}}(X^{\Delta^{1}},Y)\simeq\int_{c}Fun(X(L(c))^{\Delta^{1}},Y(L(c)))\rightarrow\int_{c}Fun(X(L(c)),Y(L(c)))^{\Delta^{1}}\simeq\\ \int_{c}Fun(X(L(c)),Y(L(c))^{\Delta^{1}})\simeq Fun_{\mathcal{B}}(X,Y^{\Delta^{1}})\simeq Fun_{\mathcal{B}}(X,Y)^{\Delta^{1}}$
@@
Note that this map is natural with respect to maps $Cat(\mathcal{B})\leftrightarrows Cat(\mathcal{A})$ induced by geometric morphisms of $\infty$-topoi.
@@
Let $Fin_{\mathcal{B}}$ be the collection of $\mathcal{B}$-categories which can be identified with locally constant functors of finite $\infty$-categories on $\mathcal{B}$. A $\mathcal{B}$-category $J$ is said to be filtered if the colimit functor $colim:\underline{Fun}_{\mathcal{B}}(I,\Omega)\rightarrow\Omega$ is continuous for each $I\in Fin_{\mathcal{B}}$.
We say that a functor $f:C\rightarrow\Omega_{\mathcal{B}}$ is filtered if it is contained in the free $Fin_{\mathcal{B}}$ cocompletion of $C$.
@@
$Filt_{\mathcal{B}}(X^{\Delta^{1}})\rightarrow Filt_{\mathcal{B}}(X)^{\Delta^{1}}.$
@@
(6) ${Fun_{\mathcal{B}}^{cc}(\underline{Psh}_{\mathcal{B}}(C),\mathcal{X})^{\Delta^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{((h_{C})^{*})^{\Delta^{1}}}$${Fun_{\mathcal{B}}(C,\mathcal{X}))^{\Delta^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}C,\Omega_{\mathcal{X}})^{\Delta^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}C,\Omega_{\mathcal{X}})^{\Delta^{1}}}$${Fun_{\mathcal{B}}^{cc}(\underline{Psh}_{\mathcal{B}}(C^{\Delta^{1}}),\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{((h_{C})^{*})}$${Fun_{\mathcal{B}}(C^{\Delta^{1}},\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Fun_{\mathcal{X}}(f^{*}(C^{\Delta^{1}}),\Omega_{\mathcal{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces Fun_{\mathcal{X}}(f^{*}(C)^{\Delta^{1}}),\Omega_{\mathcal{X}})}$
where $h_{C}$ is the internal version of the Yoneda embedding from [MW25], and the vertical maps come from the natural map of 5.4. The diagram commutes by the naturality of the construction of 5.4, Lemma 5.2 as well as the identification $f_{*}\Omega_{\mathcal{X}}=\mathcal{X}$.
@@
In this case 5.4 reduces to proving that the natural map $Filt(C(c)^{\Delta^{1}},\mathcal{X}(c))\rightarrow Filt(C(c),\mathcal{X}(c))^{\Delta^{1}}$ from 4.5 is an equivalence. Since $\mathcal{X}(c)$ is a full subcategory of $\mathcal{S}$ by Example 5.3, the result follows from Theorem 4.2.
@@
By the dual of Corollary 2.7, it suffices to show that $TOP_{\mathcal{B}}$ has a tensoring with $\Delta^{1}$ and its underlying $\infty$-category has finite colimits. The second point, is [MW25, Proposition 3.2.6.6]. For the first point, note that if $\mathcal{X}=Psh_{\mathcal{B}}(C)$, then $\mathcal{X}\otimes\Delta^{1}$ exists by Lemma 5.7. In general, choose an accessible localization $L:sPsh_{\mathcal{B}}(C)\leftrightarrows\mathcal{X}:i$. In this case, it is easy to see that if we form the pushout
@@
then $P$ satisfies the universal property of $\mathcal{X}\otimes\Delta^{1}$. ∎